Construct a​ 90% confidence interval to estimate the population mean using the data below.     x overbarequals80 sigmaequals20 nequals30 Nequals300 The​ 90% confidence interval for the population mean is left parenthesis nothing comma nothing right parenthesis . ​(Round to two decimal places as​ needed.)
Answer #1

sample mean, xbar = 80
sample standard deviation, σ = 20
sample size, n = 30


Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, Zc = Z(α/2) = 1.64


ME = zc * σ/sqrt(n)
ME = 1.64 * 20/sqrt(30)
ME = 5.99

CI = (xbar - Zc * s/sqrt(n) , xbar + Zc * s/sqrt(n))
CI = (80 - 1.64 * 20/sqrt(30) , 80 + 1.64 * 20/sqrt(30))
CI = (74.01 , 85.99)